OpenStudy (natayah):

suppose you choose a tile at random from a bag containing 5 X's, 4 Y's and 3 Z's. you replace the first tile in the bag and choose again. what is the probability of choosing 2 Y's

jimthompson5910 (jim_thompson5910):

What is the probability of choosing one Y? This will help us get toward the answer

OpenStudy (natayah):

1/12?

jimthompson5910 (jim_thompson5910):

there are 4 "y" tiles out of 5+4+3 = 9+3 = 12 total what does 4/12 reduce to?

jimthompson5910 (jim_thompson5910):

hint: divide both parts by the GCF

OpenStudy (natayah):

1/3?

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

probability of choosing one y tile = 1/3 probability of choosing two y tiles = 1/3*1/3 = 1/9 notice how I multiplied the two 1/3 terms to get 1/9

jimthompson5910 (jim_thompson5910):

This only works because we put the tile we picked back into the pile

OpenStudy (natayah):

yes! thank you, that makes a lot of sense.

jimthompson5910 (jim_thompson5910):

no problem

OpenStudy (natayah):

can you help with another problem, please?

jimthompson5910 (jim_thompson5910):

Sure. Just close this current problem and open a new thread. Post only one problem per post please.